# The Additive Bachelier model with an application to the oil option market in the Covid period

### v2

@@ Abstract
The proposed model supports efficient pricing of path-dependent exotic options via Monte Carlo simulation, using a straightforward and computationally efficient approach. Its calibration process can follow a cascade calibration: first, it accurately replicates the term structures of forwards and At-The-Money volatilities observed in the market; second, it fits the smile of the volatility surface.
@@ 1 Introduction
There exist not many models that, on the one hand, account for negative prices in oil futures and, on the other, grant a coherent description for the whole volatility surface. The Bachelier model (Bachelier 1900, Davis and Etheridge 2006) allows for negative futures prices and reproduces the term structure of At-The-Money (ATM) volatilities without considering any smile. The Bachelier model assumes that underlying prices follow a normal distribution, a significant departure from later models like Black and Scholes (1973), which assumes a log-normal distribution. Bachelier’s framework has regained relevance in modern financial contexts, particularly in markets where asset prices can take near-zero or negative values, such as interest rate derivatives and certain commodity futures. Among possible extensions, one relevant enhancement replaces standard Brownian motions with Additive processes (for a comprehensive description of the properties of these processes see, e.g., Sato 1999). These processes, which incorporate jumps and heavy-tailed distributions, improve the model’s flexibility, enabling it to capture more accurately the real-world behaviour of asset prices, especially in volatile markets. Lévy models –that require also stationarity– are the most known class among Additive processes (see, e.g., Benth et al. 2008 and Cont and Tankov 2003): unfortunately they do not have enough flexibility to offer an adequate description of some commodity markets especially in high volatile periods.
The proposed model extends the Bachelier model to a large class of Additive processes, maintaining simple tractability and a coherent description for the whole volatility surface. In particular, the proposed description offers a very simple calibration technique, known as cascade calibration: it reproduces first the most liquid derivatives, considering then the less liquid ones. It was first introduced in the interest rates market (see, e.g., Brigo and Morini 2005), where several financial products with different liquidity are involved. This elementary calibration scheme can be applied, when modelling oil derivatives even in the very volatile market of the first period of the Covid-19 pandemic, enabling to reproduce exactly both the term structure of futures and ATM volatility: this is a main result of this study.
@@
Several are the contributions of this paper. Let us briefly summarize the three most relevant ones. First, we introduce a new model in the Bachelier framework that is particularly easy to handle analytically. Second, we show that –thanks to a separability property of the Implied vol– it is possible to implement a cascade calibration, i.e., the model is calibrated first on the most liquid products (discount factors, ATM forwards), then on the ATM options, concluding with the rest of the volatility surface. Finally, we show an application to the West Texas Intermediate (WTI) oil market for the whole time window when the CME has considered Bachelier as the reference model (some months during the Covid-19 pandemic, when the market was very volatile) obtaining excellent results.
@@ 2 The model
In this Section, we present the proposed model with a detailed description of its several interesting properties. The model presents an elementary closed formula in the Fourier space for plain vanilla options (Bachelier-Lewis Formula) thanks to its explicit characteristic function. Moreover, the proposed model grants a simple description in terms of Implied volatility and admits a cascade calibration, i.e., a calibration first on the most liquid financial products and then on the less liquid ones, a key property of a model among practitioners (see, e.g., Brigo and Morini 2005).
Let us call $F_{s}$ the forward valued at time $s$ and expiry at time $t$.
We model the forward with
@@
where $\{f_{s}\}_{s\geq 0}$ a martingale process within a class of Additive processes. We call this dynamics for the forward an Additive Bachelier process, because it is a natural extension of the Bachelier model.
We model the forward (1) specifying the characteristic function for the process. The underlying process $\{f_{s}\}_{s\geq 0}$ has characteristic function exponent
@@
with $\eta\in\mathbb{R}$, $k\in\mathbb{R}^{+}$ and ${\sigma}_{s}$ a positive continuous function of time s.t. i) ${\sigma}_{s}^{2}\,s$ is increasing in time and ii) $\sigma_{s}^{2}\,s$ goes to zero for $s\rightarrow 0^{+}$. We indicate with $\mathbb{E}_{0}[\bullet]$ the expectation given the information at the value date $t_{0}=0$, with $\mathbf{p}$ the set of three model parameters: $\sigma_{t}$, $\eta$ and kk. [^1] The definition (2) explains the reason why, we require that the volatility function $\sigma_{s}$ is strictly positive. The case $\sigma_{s}=0$ for all $s\in[0,t]$ describes the trivial (and non financial) situation with the characteristic function identically equal to $1$, i.e., the forward $F_{s}$ identically equal to its initial value $F_{0}$.
The quantity $\alpha\in[0,1)$ is a hyper-parameter; it selects the chosen model: e.g. $\alpha=0$ is related to the VG model (Madan and Seneta 1990) while $\alpha=1/2$ to the NIG (Barndorff-Nielsen 1997). The function $\psi(u;k,\alpha)$ is defined as
$\psi(u;k,\alpha):=\begin{cases}\displaystyle\frac{1}{k}\displaystyle\frac{1-\alpha}{\alpha}\left\{1-\left(1+\frac{u\;k}{1-\alpha}\right)^{\alpha}\right\}&\mbox{if }\;0<\alpha<1\\[11.38109pt] \displaystyle-\frac{1}{k}\ln\left(1+u\;k\right)&\mbox{if }\;\alpha=0\;\;.\end{cases}$ (3)
Let us anticipate that the characteristic function in (2) is analytic in the horizontal strip with $\Im(u)\in(-p_{s}^{-},p_{s}^{+})$ with $p_{s}^{\pm}>0$ and finite. [^2] This result is proven in Proposition 2.10.
The Bachelier model falls within this class (as shown in Appendix A); in this case, $f_{s}$ is modelled via a Brownian motion $W_{s}$. Let us consider the European call option with strike $K$ and maturity $t$; let us call $B_{0}$ the discount factor between value date $t_{0}=0$ and maturity $t$. Within the Bachelier model, the price for the European call with moneyness $x:=K-F_{0}$ is
@@
with ${\varphi}(\bullet)$ and ${\Phi}(\bullet)$, respectively, the pdf and the CDF of a standard normal rv. The main properties of this formula are summarized in Appendix A.
The following Proposition states that the forward (modelled via (1) and the characteristic function (2)) is an Additive process (Sato 1999).
@@
Proof.
See Appendix B ∎
This is a main property of the proposed model, because –being Additive– it is possible to simulate this process with an algorithm as fast as the simple Bachelier model that requires the simulation of only Gaussian rvs (see, e.g., Azzone and Baviera 2023). Thus, any structured product with a discretely monitoring payoff can be priced and managed as a Bachelier model with similar characteristics in terms of simplicity and speed.
Moreover, as shown in the next Section, this model provides a simple closed formula for European options, which is a generalization of the Lewis (2001) formula (see, e.g., Lee 2004b).
@@ 2.1 European call price closed formula
When option’s underlying is described by (1) and $\left\{f_{t}\right\}_{t\geq 0}$ is an Additive martingale process with characteristic function $\phi_{t}$, analytic in the horizontal strip $(-p_{t}^{-},p_{t}^{+})$ with $p_{t}^{\pm}>0$, the call option can be written as
@@
for any $a\in(-p_{t}^{-},p_{t}^{+})$ and
$R_{a}=\left\{\begin{array}[]{cll}0&\text{if}&a\in(-p_{t}^{-},0)\\[8.53581pt] \displaystyle-\frac{x}{2}&\text{if}&a=0\\[8.53581pt] -x&\text{if}&a\in(0,p_{t}^{+})\end{array}\right.\;$.
Proof.
See Appendix B ∎
This closed formula for the European call allows us to price plain vanilla options in a simple and fast way, thanks to its Fourier transform formulation. Let us observe that the formula (7) is generally used in practice for $a\neq 0$ (even if it is well defined for any $a$ in the analyticity strip $(-p_{t}^{-},p_{t}^{+})$). In this case, the integrand is always analytical: $\phi_{t}$ is evaluated within its strip of analyticity and the remaining part does not contain any singularity. This closed formula can be integrated with standard quadrature techniques and even with fast Fourier transform methods (see, e.g., Press et al. 1986) that can be extremely useful speeding the calibration. The Additive Bachelier model also grants another call option formula.
@@
Proof.
See Appendix B ∎
This alternative pricing formula looks rather similar to that of the Bachelier model: the Bachelier formula (4) corresponds to the case when $G$ is identically equal to $1$, i.e. $\psi(u;k,\alpha)=-u$ that corresponds to the limiting case $k\to 0^{+}$. On the one hand, this formulation enables some important properties of the Implied volatility, as discussed in the next Section, and, on the other hand, it leads to the cascade calibration. The latter property – probably the most important characteristic of the Additive Bachelier model – is presented in Section 3.
@@ 2.2 Implied volatility and its properties
Given $\{f_{t}\}_{t\geq 0}$, $\forall{t}\in\mathbb{R}^{+}\setminus\{0\}$ and $\forall{x}\in\mathbb{R}$, it exists a unique ${\cal I}_{t}(x)\in\mathbb{R}^{+}$ solution of the IV equation (9).
Proof.
See Appendix B ∎
@@
The Implied vol ${\cal I}_{t}(x)$ is a separable function of $y$ (10) and $t$
${\cal I}_{t}(x)=\sigma_{t}\,I(y)\;,$ (11)
where I(y){I}(y) is time independent and can be obtained by solving the following equation
${c_{b}\left(y,{I}(y)\right)}=\mathbb{E}\left[c_{b}\left(y+\eta\,(G-1),\sqrt{G}\right)\right]\;.$ (12)
Proof.
See Appendix B ∎
By abuse of language, whenever there is no ambiguity, we also call $I(y)$ Implied vol or volatility smile, since it describes the volatility smile of the Implied vol.
@@
Proof.
See Appendix B ∎
Remark. It is possible to observe that for $k$ equal to zero, the unique solution of the IV equation (12) is I(y)=1{I}({y})=1 $\forall{y}\in\mathbb{R}$, corresponding to the Bachelier model with time dependent deterministic volatility.
@@
I{I} is an even function of y{y}, iff $\eta=0$.
Proof.
See Appendix B ∎
The above properties suggest a comment on the parsimony of the proposed model. Not only the Additive Bachelier has few parameters –just three– being able to model the volatility surface. They also describe three main distinct aspects of the surface, as we’ll discuss in detail in Section 3, $\sigma_{t}$ models the term structure of the ATM volatility, i.e. it controls the level of the volatility surface. Proposition 2.7 and Corollary 2.8 state that $\eta$ models the skew, controlling the symmetry of the volatility, and for $\eta=0$ describes the surface with a perfectly symmetric smile around the ATM. Finally, $k$ models the vol-of-vol, i.e. the convexity of the surface, being the limit $k\to 0^{+}$ the classical Bachelier model with no smile.
@@
The Implied volatility I{I} around the ATM is
@@
$\left\{\begin{array}[]{ccrl}{I}_{0}&=&\sqrt{2\pi}&\mathbb{E}\left[c_{b}\left({\eta}\,(G-1),\sqrt{G}\right)\right]\\[14.22636pt] {I}^{\prime}_{0}&=&\displaystyle-\sqrt{\frac{\pi}{2}}&\displaystyle\mathbb{E}\left[{\rm erf}\left(\frac{{\eta}}{\sqrt{2}}\,\frac{1-G}{\sqrt{G}}\right)\right]\\[14.22636pt] {I}^{\prime\prime}_{0}&=&\displaystyle\sqrt{2\pi}&\displaystyle\mathbb{E}\left[\frac{1}{\sqrt{G}}\,\varphi\left({\eta}\,\frac{1-G}{\sqrt{G}}\right)\right]-\frac{1}{I_{0}}\end{array}\right$.\;, (13)
with $I^{\prime}_{0}$ an odd function in $\eta$.
Proof.
See Appendix B ∎
These results are extremely relevant for both ATM volatility and volatility skew.
@@
$\sigma^{ATM}_{t}:=\mathscr{I}_{t}(x=0)=\sigma_{t}\,I(y=0)=\sigma_{t}\,I_{0}$ (14)
i.e. the ATM volatility is proportional to $\sigma_{t}$, with the proportionality coefficient equal to ${I}_{0}$ in (13) that depends only on $\eta$ and $k$. Moreover, the quantity ${I}_{0}$ is strictly positive due to the Jensen inequality
$I_{0}=\sqrt{2\,\pi}\,\mathbb{E}\left[-(\eta\,(G-1)+\sqrt{G}g)\right]^{+}>\sqrt{2\,\pi}\,\left[\mathbb{E}(-\eta\,(G-1)-\sqrt{G}g)\right]^{+}=0\;\;.$
@@
The following result extends Lee (2004a) formula, providing the characterisation of the asymptotic behaviour of the Implied vol in the Additive Bachelier model. Let us notice that the asymptotic behaviour is considered for the limit in Proposition 2.10, while Lee (2004a) considers the $\lim\sup$.
@@
The asymptotic behaviour of I(y){I}({y}) for large |y||{y}| is
$\lim_{{y}\to\pm\infty}\frac{{I}({y})^{2}}{|{y}|}=\frac{1}{2\;p^{\mp}}$
@@
$p^{\pm}=\mp\eta+\sqrt{{\eta}^{2}+2\,\frac{1-\alpha}{k}}\>.$
Proof.
See Appendix B ∎

### v3

@@ Abstract
Model calibration can follow a cascade procedure: first, it accurately replicates the term structures of forwards and At-The-Money volatilities observed in the market; second, it fits the smile of the volatility surface. The proposed model also supports efficient pricing of path-dependent exotic options via Monte Carlo simulation, using a straightforward and computationally efficient approach.
@@ 1 Introduction
There exist not many models that, on the one hand, account for negative prices in oil futures and, on the other, grant a coherent description for the whole volatility surface. The Bachelier model (Bachelier 1900, Davis and Etheridge 2006) allows for negative futures prices and reproduces the term structure of At-The-Money (ATM) volatilities without considering any smile. The Bachelier model assumes that underlying prices follow a normal distribution, a significant departure from later models like Black and Scholes (1973), which assumes a log-normal distribution. Bachelier’s framework has regained relevance in modern financial contexts, particularly in markets where asset prices can take near-zero or negative values, such as interest rate derivatives and certain commodity futures. Among possible extensions, one relevant enhancement replaces standard Brownian motions with additive processes (for a comprehensive description of the properties of these processes see, e.g., Sato 1999). These processes, which incorporate jumps and heavy-tailed distributions, improve the model’s flexibility, enabling it to capture more accurately the real-world behaviour of asset prices, especially in volatile markets. Lévy models –that require also stationarity– are the most known class among additive processes (see, e.g., Benth et al. 2008 and Cont and Tankov 2003): unfortunately, they do not have enough flexibility to offer an adequate description of some commodity markets, especially in highly volatile periods.
The proposed model extends the Bachelier model to a large class of additive processes, maintaining simple tractability and a coherent description for the whole volatility surface. In particular, the proposed description offers a very simple calibration procedure, known as cascade calibration: it reproduces first the most liquid derivatives, considering then the less liquid ones. This technique was first introduced in the interest rates market (see, e.g., Brigo and Morini 2005), where several financial products with different liquidity are involved. This elementary calibration scheme can be applied when modelling oil derivatives, even in the very volatile market of the first period of the Covid-19 pandemic, enabling to reproduce exactly both the term structure of futures and ATM volatility: this is a main result of this study.
@@
Several are the contributions of this paper. Let us briefly summarize the three most relevant ones. First, we introduce a new model in the Bachelier framework that is particularly easy to handle analytically. Second, we show that –thanks to a separability property of the implied vol– it is possible to implement a cascade calibration, i.e. the model is calibrated first on the most liquid products (discount factors, ATM forwards), then on the ATM options, concluding with the rest of the volatility surface. Finally, we show an application to the West Texas Intermediate (WTI) oil market for the whole time window when the CME has considered Bachelier as the reference model (some months during the Covid-19 pandemic, when the market was very volatile), obtaining excellent results.
@@ 2 The model
In this Section, we present the proposed model with a detailed description of its several interesting properties. The model presents an elementary closed formula in the Fourier space for plain vanilla options (Bachelier-Lewis Formula) thanks to its explicit characteristic function. Moreover, the proposed model grants a simple description in terms of implied volatility and admits a cascade calibration, i.e. a calibration first on the most liquid financial products and then on the less liquid ones, a key property of a model among practitioners (see, e.g., Brigo and Morini 2005).
Let us consider a filtered probability space $(\Omega,\mathscr{F},\mathbb{F},\mathbb{P})$, where the filtration $\mathbb{F}=(\mathscr{F}_{s})_{s\geq 0}$ is right-continuous and complete. Let us call $F_{s}$ the forward valued at time $s$ and expiry at time $t$.
We can model the forward with
@@
where $\{f_{s}\}_{s\geq 0}$ is an $\mathbb{F}$-adapted process, with $f_{0}=0$, within a class of additive processes, defined by specifying their characteristic exponent as follows
$\ln\phi_{s}(u):=\ln\mathbb{E}_{0}\left[e^{i\,u\,f_{s}}\right]=\psi\left(i\,u\,\eta_{s}\,{\sigma}_{s}\,\sqrt{s}+\frac{u^{2}}{2}\,{\sigma}_{s}^{2}\,s;\;k_{s},\;\alpha\right)\,+\,i\,u\,\eta_{s}\,{\sigma}_{s}\,\sqrt{s}\;\;,$ (2)
with $\sigma_{s},\,\eta_{s},\,k_{s}$ continuous function of time. We indicate with $\mathbb{E}_{0}[\bullet]$ the expectation given the information at the value date $t_{0}=0$.
The function $\psi(u;k,\alpha)$ is defined as
$\psi(u;k,\alpha):=\begin{cases}\displaystyle\frac{1}{k}\displaystyle\frac{1-\alpha}{\alpha}\left\{1-\left(1+\frac{u\;k}{1-\alpha}\right)^{\alpha}\right\}&\mbox{if }\;0<\alpha<1\\[11.38109pt] \displaystyle-\frac{1}{k}\ln\left(1+u\;k\right)&\mbox{if }\;\alpha=0\end{cases}\,\,.$ (3)
and it is the Laplace exponent of a tempered stable random variable. [^1] Although this characteristic exponent resembles that of a subordinated process, it is useful to emphasize that $\{f_{s}\}_{s\geq 0}$ cannot be obtained in general via subordination. The only exception is the specific case where it reduces to a Lévy process. A similar example is discussed in Azzone and Baviera (2025). The quantity $\alpha\in[0,1)$ is a hyper-parameter; it selects the chosen model: e.g. $\alpha=0$ is related to the VG model (Madan and Seneta 1990) while $\alpha=1/2$ to the NIG (Barndorff-Nielsen 1997).
An additive process is defined as a càdlàg adapted stochastic process on $\mathbb{R}$, denoted by $\{f_{s}\}_{s\geq 0}$ with $f_{0}=0$ $\mathbb{P}$-a.s., which is uniquely characterized by the property of independent increments and stochastic continuity (see, e.g., Cont and Tankov 2003, Def.14.1). According to Theorem 9.1 in Sato (1999), the distribution of an additive process at any time $s$ is infinitely divisible, and it is identified by the generating triplet $(A_{s},\gamma_{s},\nu_{s})$ (see, e.g., Sato 1999, Def.8.2, p.38). The fundamental relationship between a system of infinitely divisible distributions and the existence of an underlying additive process is established by Theorem 9.8 in Sato (1999). This Theorem prescribes two essential sets of conditions on the generating triplet: firstly, monotonicity requirements to ensure the non-negativity of diffusion terms and Lévy measures for the increments; and secondly, continuity conditions to guarantee the stochastic continuity of the process. By invoking these results, we demonstrate in Appendix Appendix : Bachelier formula, Lévy Bachelier and its generalization that there exists a family of additive processes uniquely defined by the characteristic exponent (2).
In this paper, we model the forward (1) considering a sub-case of the processes in (2). The underlying process $\{f_{s}\}_{s\geq 0}$ has characteristic function exponent
@@
with $\eta\in\mathbb{R}$, $k\in\mathbb{R}^{+}$ and ${\sigma}_{s}$ a positive continuous function of time s.t. i) ${\sigma}_{s}^{2}\,s$ is increasing in time and ii) $\sigma_{s}^{2}\,s$ goes to zero for $s\rightarrow 0^{+}$. [^2] The definition (4) explains the reason why, we require that the volatility function $\sigma_{s}$ is strictly positive. The case $\sigma_{s}=0$ for all $s\in[0,t]$ describes the trivial (and non financial) situation with the characteristic function identically equal to $1$, i.e. the forward $F_{s}$ identically equal to its initial value $F_{0}$. Let us emphasise that additive processes with characteristic function (2) include the Lévy processes as the stationary sub-case, as it is shown in Appendix Appendix : Bachelier formula, Lévy Bachelier and its generalization. Our model is not included in the Lévy sub-case. We thank one reviewer for addressing this point. We indicate with $\mathbf{p}$ the set of three model parameters: $\sigma_{s}$, $\eta$ and kk. [^3] We would indicate the characteristic function (4) with $\phi_{s}(u;\mathbf{p})$, but we keep the dependence on $\mathbf{p}$ implicit when not required for clarity.
Let us anticipate that the characteristic function in (4) is analytic in the horizontal strip with $\Im(u)\in(-p^{+}_{s},p^{-}_{s})$ with $p^{\pm}_{s}>0$ and finite. [^4] This result is proven in Proposition 2.10.
We call this dynamics for the forward an additive Bachelier process, because it is a natural extension of the Bachelier model.
The Bachelier model falls within this class (as shown in Appendix Appendix : Bachelier formula, Lévy Bachelier and its generalization); in this case, $f_{s}$ is modelled via a Brownian motion $W_{s}$. Let us consider the European call option with strike $K$ and maturity $t$; let us call $B_{0}$ the discount factor between value date $t_{0}=0$ and maturity $t$. Within the Bachelier model, the price for the European call with moneyness $x:=K-F_{0}$ is
@@
with ${\varphi}(\bullet)$ and ${\Phi}(\bullet)$, respectively, the pdf and the cdf of a standard normal rv. The main properties of this formula are summarised in Appendix Appendix : Bachelier formula, Lévy Bachelier and its generalization.
The following Proposition states that the forward (modelled via (1) and the characteristic function (4)) is an additive and martingale process. The proof of the following Proposition, along with all other proofs in this paper, is provided in Appendix Appendix : Proofs.
@@
This property of the proposed model is important: being additive, it is possible to simulate this process with an algorithm as fast as the simple Bachelier model that requires the simulation of only Gaussian rvs (see, e.g., Azzone and Baviera 2023). Thus, any structured product with a discretely monitored payoff can be priced and managed as a Bachelier model with similar characteristics in terms of simplicity and speed.
Moreover, as shown in the next Section, this model provides a simple closed formula for European options, which is a generalization of the Lewis (2001) formula.
@@ 2.1 European call price closed formula
When option’s underlying is described by (1) and $\left\{f_{t}\right\}_{t\geq 0}$ is an additive martingale process with characteristic function $\phi_{t}$, analytic in the horizontal strip $(-p^{+}_{t},p^{-}_{t})$ with $p^{\pm}_{t}>0$, the call option can be written as
@@
for any $a\in(-p^{+}_{t},p^{-}_{t})$ and
$R_{a}=\left\{\begin{array}[]{cll}0&\text{if}&a\in(-p^{+}_{t},0)\\[8.53581pt] \displaystyle-\frac{x}{2}&\text{if}&a=0\\[8.53581pt] -x&\text{if}&a\in(0,p^{-}_{t})\end{array}\right.\;$.
This closed formula for the European call allows us to price plain vanilla options in a simple and fast way, thanks to its Fourier transform formulation. Let us observe that the formula (8) is generally used in practice for $a\neq 0$ in the analyticity strip $(-p^{+}_{t},p^{-}_{t})$. In this case, the integrand is always analytic: $\phi_{t}$ is evaluated within its strip of analyticity, and the remaining part does not contain any singularity. This closed formula can be integrated with standard quadrature techniques and even with fast Fourier transform methods (see, e.g., Press et al. 1986) that can be extremely useful, speeding up the calibration. The additive Bachelier model also grants another call option formula.
@@
This alternative pricing formula looks rather similar to that of the Bachelier model. [^5] The integral in (9) is computed in the real space for $\alpha=0,1/2$, when the pdf of $G$ is known explicitly (see, e.g., Cont and Tankov 2003, Table 4.4, Sec.4.4.2), while for the other values of $\alpha$ is computed in the Fourier space with a formula like the Bachelier-Lewis (8). The Bachelier formula (5) corresponds to the case when the rv $G$ is identically equal to $1$, i.e. $\psi(u;k,\alpha)=-u$ that corresponds to the limiting case $k\to 0^{+}$. On the one hand, this formulation enables some important properties of the implied volatility, as discussed in the next Section, and on the other hand, it leads to the cascade calibration. The latter property – probably the most important characteristic of the additive Bachelier model – is presented in Section 3.
@@ 2.2 Implied volatility and its properties
Given $\{f_{t}\}_{t\geq 0}$, for any ${t}\in\mathbb{R}^{+}\setminus\{0\}$ and ${x}\in\mathbb{R}$, it exists a unique ${\cal I}_{t}(x)\in\mathbb{R}^{+}$ solution of the IV equation (10).
@@
For any $y\in\mathbb{R}$, $t\in\mathbb{R}^{+}\setminus\{0\}$, the implied vol ${\cal I}_{t}(\bullet)$ is a separable function of $y$ (11) and $t$
${\cal I}_{t}\left(y\,\sigma_{t}\,\sqrt{t}\right)=\sigma_{t}\,I(y)\;,$ (12)
where ${I}(\bullet)$ is the solution of the time independent equation
${c_{b}\left(y,{I}(y)\right)}=\mathbb{E}\left[c_{b}\left(y+\eta\,(G-1),\sqrt{G}\right)\right]\;,$ (13)
where $G$ is the rv in (9).
Let us notice that the above property is a fundamental characteristic of this model. The possibility to separate the dependency due to the maturity in the implied volatility will allow us to develop a calibration scheme that reproduces exactly ATM vols as described in Section 3. By abuse of language, whenever there is no ambiguity, we also call $I(y)$ implied vol or volatility smile, since it describes the volatility smile of the implied vol. [^6] As noted above regarding the characteristic function $\phi_{s}(u)$, the implied vol would be formally denoted by $I(y;\eta,k)$. However, to avoid cluttering the notation, we write $I(y)$, keeping the dependence on $\eta$ and $k$ implicit when it is clear from the context. In instances where only one of the two parameters is relevant, e.g. $\eta$, we denote the volatility smile as $I(y;\eta)$.
@@
It is possible to observe that for $k$ equal to zero, the unique solution of the IV equation (13) is I(y)=1{I}({y})=1 $\forall{y}\in\mathbb{R}$, corresponding to the Bachelier model with time dependent deterministic volatility in (5).
@@
For any $k\in\mathbb{R}^{+}$, ${I}(y;\eta)$ is an even function of y{y}, iff $\eta=0$.
The above properties suggest a comment on the parsimony of the proposed model. Not only does the additive Bachelier have few parameters –just three– being able to model the volatility surface. They also describe three main distinct aspects of the surface, as we’ll discuss in detail in Section 3, $\sigma_{t}$ models the term structure of the ATM volatility, i.e. it controls the level of the volatility surface. Proposition 2.7 and Corollary 2.8 state that $\eta$ models the volatility skew, controlling the symmetry of the volatility, and for $\eta=0$ describes the surface with a perfectly symmetric smile wrt the ATM. Finally, $k$ models the vol-of-vol, i.e. the convexity of the surface, being the limit $k\to 0^{+}$ the classical Bachelier model with no smile.
@@
For any $\eta\in\mathbb{R}$, $k\in\mathbb{R}^{+}$, the implied volatility I{I} around the ATM (i.e., for $y\to 0$) is
@@
$\left\{\begin{array}[]{ccrl}{I}_{0}:=&I(0)&=&\sqrt{2\pi}\;\,\mathbb{E}\left[c_{b}\left({\eta}\,(G-1),\sqrt{G}\right)\right]\\[14.22636pt] {I}^{\prime}_{0}:=&\displaystyle\left.\frac{\partial{{I}}(y)}{\partial y}\right|_{y=0}&=&\displaystyle-\sqrt{\frac{\pi}{2}}\displaystyle\;\,\mathbb{E}\left[{\rm erf}\left(\frac{{\eta}}{\sqrt{2}}\,\frac{1-G}{\sqrt{G}}\right)\right]\\[14.22636pt] {I}^{\prime\prime}_{0}:=&\displaystyle\left.\frac{\partial^{2}{{I}}(y)}{\partial y^{2}}\right|_{y=0}&=&\displaystyle\sqrt{2\pi}\;\displaystyle\;\,\mathbb{E}\left[\frac{1}{\sqrt{G}}\,\varphi\left({\eta}\,\frac{1-G}{\sqrt{G}}\right)\right]-\frac{1}{I_{0}}\end{array}\right$.\;, (14)
with $I^{\prime}_{0}$ an odd function in $\eta$ and $G$ is a positive rv with Laplace exponent $\psi(u;k,\alpha)$, defined in (3).
These results are extremely relevant for both ATM volatility and volatility skew. [^7] The integrals in (14) are computed in the real space for $\alpha=0,1/2$, as in footnote 5.
@@
$\sigma^{ATM}_{t}:=\mathscr{I}_{t}(x=0)=\sigma_{t}\,I_{0}$ (15)
i.e. the ATM volatility is proportional to $\sigma_{t}$, with the proportionality coefficient equal to ${I}_{0}$ in (14) that depends only on $\eta$ and $k$. Moreover, the quantity ${I}_{0}$ is strictly positive, because it is the expectation of a positive function.
@@
The following result extends Lee (2004b) formula, providing the characterisation of the asymptotic behaviour of the implied vol in the additive Bachelier model. Let us notice that the asymptotic behaviour is considered for the limit in Proposition 2.10, while Lee (2004b) considers the $\lim\sup$.
@@
For any $\eta\in\mathbb{R}$, $k\in\mathbb{R}^{+}$, the asymptotic behaviour of I(y){I}({y}) for large |y||{y}| is
$\lim_{{y}\to\pm\infty}\frac{{I}({y})^{2}}{|{y}|}=\frac{1}{2\;p^{\pm}}$
@@
$p^{\pm}=\pm\eta+\sqrt{{\eta}^{2}+2\,\frac{1-\alpha}{k}}\>.$